x + 2 / (x + 1) * (x + 2) = x - 2 / (x + 1) * (x - 2)
Answer by Boreal(15235)

x + 2 / (x + 1) * (x + 2) = x - 2 / (x + 1) * (x - 2)
multiply everything by (x+1)(x+2)(x-2)
get (x+2)(x-2)=(x-2)(x+2), which is an identity,
Perhaps it is (x+2)/(x+1) all multiplied by (x+2)=(x-2)/(x+1) all multiplied by (x-2)
Then (x^2+4x+4)/(x+1)=(x^2-4x+4)/(x+1)
The denominators are equal, so just deal with the numerators, and the x^2 cancel
4x+4=-4x+4
8x=0
x=0 ANSWER
(2/1)(2)=(-2/1)*(-2)
4=4